OpenStudy (anonymous):

Need help with quadratic equations Solve: -3x^3 - 4x - 4 = 0

OpenStudy (anonymous):

which methods are you familiar with? can you factor? use the quadratic equation? complete the square?

OpenStudy (anonymous):

I'm familiar with the quadratic equation

OpenStudy (anonymous):

so why don't you use that?

OpenStudy (anonymous):

ooh, sorry... didn't see the cube. is it \[-3x^3-4x-4=0\]

OpenStudy (anonymous):

or is it \[-3x^2-4x-4=0\]

OpenStudy (anonymous):

No its squared and when i do it i get up to 4+/-sqrt(-32)/-6 and get stuck

OpenStudy (anonymous):

the second one

OpenStudy (anonymous):

so here's the deal... "can't" take the square root of a negative number, right? well, turns out you can. we call \(i\) the imaginary number. \(i=\sqrt{-1}\). so anytime you have to take the square root of a negative number, you factor out \(i\). for example: suppose you have to take the square root of -16...\[\sqrt{-16}=\sqrt{16}\cdot \sqrt{-1} = 4\cdot i = 4i\]

OpenStudy (anonymous):

it is customary to put rational numbers (integers and fractions) before \(i\) and irrational numbers (square, cube, etc. roots) after \(i\). That's why I wrote it as \(4i\) and not \(i4\).

OpenStudy (anonymous):

oh ok, could you help me simplify it more? because its a practice test and the answers are

OpenStudy (anonymous):

http://imgur.com/RlyOSLp

OpenStudy (anonymous):

\[\text{When }ax^+bx+c=0 \text{, then }\frac{ -b \pm \sqrt{b^2-4ac} }{ 2a } \Rightarrow \frac{ -\left( -4 \right) \pm \sqrt{\left( -4 \right)^2-4\left( -3 \right) \left( -4 \right)} }{ 2\left( -3 \right) }\]\[\Rightarrow \frac{ 4 \pm \sqrt{16-48} }{ -6 }=\frac{ 4 \pm \sqrt{-32} }{ -6 }=\frac{ 4 \pm \sqrt{16\cdot2}\cdot \sqrt{-1} }{ -6 }=\frac{ 4 \pm 4i \sqrt{2} }{ -6 }=-\frac{ 2 }{3 }\pm\frac{ 2i }{3 }\]

OpenStudy (anonymous):

or as they have it...\[=\frac{ -2\pm 2i \sqrt{2} }{3 }\]

OpenStudy (anonymous):

learn to do this work yourself and you WILL get better at it. use this site to have people check your work and help with explanations.

OpenStudy (anonymous):

Oh i see what you do now, thank you so much for your help!